# AstroMLab 3: Achieving GPT-4o Level Performance in Astronomy with a Specialized 8B-Parameter Large Language Model

### v1

@@ VI Availability
The code used to prepare the datasets and perform the training will be made available upon reasonable request.
@@ VII Discussion and Future Work
This work demonstrates the potential of specialized language models in astronomy through a systematic approach to model development and evaluation. While previous efforts like laid important groundwork in domain-specific modeling, the field has faced persistent challenges in achieving performance gains over baseline models, especially in instruction-following tasks. Our multi-stage training process—combining continued pretraining, extensive supervised fine-tuning, and strategic model merging—addresses these challenges, achieving a notable improvement over the baseline model.
@@ Acknowledgements
This research used resources of the Oak Ridge Leadership Computing Facility (OLCF), which is a DOE Office of Science User Facility at the Oak Ridge National Laboratory supported by the U.S. Department of Energy under Contract No. DE-AC05-00OR22725 and support from Microsoft’s Accelerating Foundation Models Research (AFMR) program. TdH was supported by World Premier International Research Center Initiative (WPI), MEXT, Japan. YST is supported by the National Science Foundation under Grant No. 2406729.

### v2

@@ VI Availability
The code used to prepare the datasets and perform the training will be made available upon reasonable request. The synthetically generated Q&\&&A pairs are currently supporting research and development of our next-generation model. After the completion of our research trajectory, we anticipate a full release of our synthetic data set. For inquiries regarding data usage, collaboration, or access to specific subsets of our work, interested parties are encouraged to contact the corresponding author.
@@ VII Discussion and Future Work
This work demonstrates the potential of specialized language models in astronomy through a systematic approach to model development and evaluation. While previous efforts like Pan et al. [7] laid important groundwork in domain-specific modeling, the field has faced persistent challenges in achieving performance gains over baseline models, especially in instruction-following tasks. Our multi-stage training process—combining continued pretraining, extensive supervised fine-tuning, and strategic model merging—addresses these challenges, achieving a notable improvement over the baseline model.
@@ Acknowledgements
This research used resources of the Oak Ridge Leadership Computing Facility (OLCF), which is a DOE Office of Science User Facility at the Oak Ridge National Laboratory supported by the U.S. Department of Energy under Contract No. DE-AC05-00OR22725 and support from Microsoft’s Accelerating Foundation Models Research (AFMR) program. TdH was supported by World Premier International Research Center Initiative (WPI), MEXT, Japan. YST is supported by the National Science Foundation under Grant No. 2406729. Work at Argonne National Lab is supported by UChicago Argonne LLC, Operator of Argonne National Laboratory. Argonne, a U.S. Department of Energy Office of Science Laboratory, is operated under contract no. DE-AC02-06CH11357. A special thanks goes out to Cassie Reuter and Joshua Montgomery for acting as independent evaluators.